• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 04:32:33 on March 4,2012

    Entry number 363340

    keyword=Left septum power off for runs 2682-2698.

    From run 2682, left arm had very low rates. Checked a bunch of things (HRS magnets settings, S1/S2m HVs, CODA reboot, etc.), turned out the left septum power was off. Turned it back on.
    Set the current back. All set and ready for data taking again.
    runs 2682 to 2698 had left septum off.